A five-year bond with a yield of 7% (continuously compounded) pays an 8% coupon at the end of each year.

a) What is the bond’s price?

b) What is the bond’s duration?

c) Use the duration to calculate the effect on the bond’s price of a 0.2% decrease in its yield.

d) Recalculate the bond’s price on the basis of a 6.8% per annum yield and verify that the result is in agreement with your answer to (c).
Answer #1

ra (a) bet loud Parce 1000 Let bond face value - 1000 cov) coup on rate a 87 (5) yield = 7) (9) Coupon - 1000 x 8) = 0 PriceDuration (6) Bond founda = (PV, 1) + (PV 2 x 2) + (P ve x 3) + (Pv4 x 4) + (PVS XS). Pr + PV2 + PV2 + PV y + PV5 PV = Presenta change in Londs luu Potion Dmac & Bond duration or Macaulays dy = change in yold y - original yold dy = 0.2% mac = 4.32 y| لنا = 329.779719.687 News and Price & 11049.466] % change in Price = 1049.466 - 1047.00) 70 4.001 X 700 which same as is in